OpenStudy (anonymous):

how do i find the two asymptotes and period for y=tan(x/3)?

OpenStudy (sidsiddhartha):

y=tan(x/3)=sin(x/3)/cos(x/3) Asymptotes occur when cosx/3=0. x=(2n+1)*(pi/2) put n=0,1,2 and generally period of tan(kx)=pi/k

OpenStudy (anonymous):

why am i putting n=0,1 & 2 if i only need 2 asymptotes?

OpenStudy (sidsiddhartha):

it should be 0,1

OpenStudy (anonymous):

so for 0 i got pi/2

OpenStudy (anonymous):

and for 1 i got 3pi/2. would those be the two asymptotes?

OpenStudy (anonymous):

or would one of the asymptotes be 3?

OpenStudy (sidsiddhartha):

it should be x/3=(2n+1)*(pi/2)

OpenStudy (anonymous):

okay so then 3pi/2 and 9pi/2 is what i got. is that correct?

OpenStudy (anonymous):

and the period would be pi/3?

OpenStudy (sidsiddhartha):

here k=1/3

OpenStudy (anonymous):

is k=1/3 the period of part of the formula to find the asymptotes?

OpenStudy (sidsiddhartha):

i think asymptotes are all right and for period p=pi/k=pi/(1/3)=3(pi)

OpenStudy (anonymous):

oh okay, thank you for your helP! :)